a sky jumper speeds down a slope and exits the ski track in horizontal direction at a velocity of 25 m per second where the landing inclined falls of the slope of 33 °

solve the following

what is the halftime of flight of the sky jumper

what is the range of the sky jumper when it lands


1
Expert's answer
2021-10-14T11:35:25-0400

Determine the height between the point of jump and point of landing:


"h=\frac{gt^2}2."

Determine the range:


"R=vt."

The relation of R to h is


"h=R\tan\theta=vt\tan\theta.\\\space\\
vt\tan\theta=\frac{gt^2}2,\\\space\\
t=\frac{2v\tan\theta}{g}=3.31\text{ s},"

the halftime is 1.66 s, the range is


"R=25·3.31=82.75\text{ m}."